# Learning to Group and Grasp Multiple Objects

### v1

@@ Abstract
Simultaneously grasping and transporting multiple objects can significantly enhance robotic work efficiency and has been a key research focus for decades. The primary challenge lies in determining how to push objects, group them, and execute simultaneous grasping for respective groups while considering object distribution and the hardware constraints of the robot. Traditional rule-based methods struggle to flexibly adapt to diverse scenarios. To address this challenge, this paper proposes an imitation learning-based approach. We collect a series of expert demonstrations through teleoperation and train a diffusion policy network, enabling the robot to dynamically generate action sequences for pushing, grouping, and grasping, thereby facilitating efficient multi-object grasping and transportation. We conducted experiments to evaluate the method under different training dataset sizes, varying object quantities, and real-world object scenarios. The results demonstrate that the proposed approach can effectively and adaptively generate multi-object grouping and grasping strategies. With the support of more training data, imitation learning is expected to be an effective approach for solving the multi-object grasping problem.
@@
Multi-Object Grasping, Manipulation, Imitation Learning
@@ I Introduction
In recent years, advances in AI and sensor technology have enabled the development of robots capable of performing picking tasks in environments where objects are randomly scattered. In such situations, robots need to transport objects efficiently to their target locations. However, a common limitation of robots is that they can typically handle only one object at a time, which leads to slower task execution. Different from robots, humans can grasp and carry multiple objects at once, and are thus more efficient. To bridge this gap, this study focuses on enabling robots to perform simultaneous multi-object grasping, thereby improving their work efficiency.
@@
Previous studies on simultaneous multi-object grasping has identified several challenges. For example, Sakamoto et al. [1] proposed a rule-based method to push objects closer together before grasping them simultaneously. However, the method cannot handle complex pushing actions, and the combinations of objects that can be grasped together are limited. Agboh et al. [2] studied identifying near objects arranged on a plane and grasping multiple objects by pushing them together as the gripper closes. The method works under specific conditions that objects are not spaced farther apart than the gripper’s maximum opening width. Or else, it will degenerate to single-object grasping. These challenges highlight the need for better strategies to group objects through complex pushing actions and to select suitable combinations of objects for simultaneous grasping. Unfortunately, making such decisions based on fixed rules is difficult.
@@
Figure 1: (a) Collecting human expert data using teleoperation. (b) The collected dataset is used to train a diffusion policy network for inferring actions. (c) The policy network uses past observations to generate appropriate pushing or grasping actions.
In this context, we propose addressing these issues by using imitation learning and enabling robots to learn how humans decide which objects to group and grasp. The goal is to allow robots to perform simultaneous multi-object grasping in scattered environments. By acquiring grasping strategies through imitation learning, the proposed method enables robots to automatically choose suitable object combinations and perform the necessary pushing actions to group them. The method can overcome the limitations of existing methods and improve the efficiency and versatility of robotic picking tasks.
@@
Specifically, the proposed framework comprises three main components: data collection, learning, and inference. Fig. 1 illustrates these components and workflow. In the data collection phase shown in Fig. 1(a), a teleoperation system is developed by integrating a vision-based hand-detection model, allowing human operators to control the robot using natural hand movements. The setup enables collecting data like robot joint angles, gripper opening and closing states, and images from by the global and local cameras mounted in the environment and on the robot hand. In the learning phase shown in Fig. 1(b), the framework employs Diffusion Policy [3] to learn appropriate robot actions several steps ahead, based on the robot’s state from previous steps. Finally, in the inference phase shown in Fig. 1(c), the framework uses the robot’s joint angles, gripper states, and camera images observed from a few steps before to predict joint angles and gripper states for subsequent steps, thereby generating simultaneous grasping motions for multiple objects. The trained neural network learns policies from human demonstrations to perform object grouping and grasping, thus significantly improving robotic picking efficiency.
@@
Two groups of experiments were conducted to validate the proposed framework. The first group focused on comparing and analyzing the influence of different training data volumes and object quantities on learning and inference. The second group conducted fine-tuning a model pre-trained on a dataset of simple objects to more practical scenarios and evaluated the generality of the proposed method to real-world objects. The experimental results validated the effectiveness of the proposed method and validated the adaptivity of fine-tuned models. Additionally, the results indicated that training a well-performing policy network requires a large amount of high-quality data, with a positive correlation between data volume and learning performance. Developing efficient methods for collecting such data could potentially improve both learning and inference performance.
@@ II Related Work
Multi-object simultaneous grasping has been attracting significant attention for over two decades. As pioneering contributions, Aiyama et al. [4] studied using two manipulators to cooperatively grasp and lift an array of box-shaped objects. Harada and Kaneko [5][6] studied the neighboring equilibrium of multiple cylindrical objects and their enveloping grasps. Yoshikawa et al. [7] proposed the optimal power grasp condition for multiple objects from the viewpoint of reducing finger joint torques. Yamada et al. [8] presented a more general analysis of the grasp stability of multiple objects grasped by fingers considering objects’ surface curvatures and the potential energy of the spring models in contact. The authors later extended the theories to multi-fingered hands with revolute joints [9] and 3D objects [10].
@@
More recent work by Chen et al. [11] inserted the Barrett hand into a pile of objects and estimated the number of grasped items using a deep learning model. Shenoy et al. [12] further focused on efficiently transferring the grasped multiple objects to a separate container. Takahashi et al. [13] applied multi-object grasping evaluation to grasping a given amount of granular foods. The authors also explored post-grasping methods for fine adjustments [14]. Li et al. [15] developed a framework for planning and learning to grasp (lift) multiple objects. Jiang et al. [16] developed a vacuum gripper equipped with multiple suction cups and also the grasp planning methods to simultaneously pick multiple objects from a tray. Nguyen et al. [17] proposed the Wiring-Claw Gripper, which took advantage of the soft interaction between wires in a claw and objects for simultaneous multi-object grasping.
@@
The above approaches achieved stable grasping of multiple or unknown objects. However, they often assumed that the objects were initially positioned in proximity and thus are less applicable to environments where objects are scattered. To tackle such limitations, several studies have explored approaches that utilize pushing actions to rearrange objects into positions more conducive to simultaneous grasping. For example, Sakamoto et al. [1] proposed a method for grasping two distant objects by first pushing one object closer to the other and then grasping them together. The approach determined whether to grasp a single object, grasp two objects simultaneously, or push one object closer for simultaneous grasping based on the distance and friction between objects. Using dynamic programming, the authors optimized the grasping sequence and enabled simultaneous grasping of two rectangular objects by rotating them about the intersection of their long axes. Agboh et al. [2][18] proposed the $\mu$-MOG framework to simultaneously grasp multiple rigid convex polygonal objects scattered on a plane. The method facilitated the simultaneous grasping of cluttered objects by aligning them using a gripper’s opening jaw. Shrey et al. [19] partitioned clusters to ensure that the sum of the minimum graspable diameters within a cluster remained below a gripper’s width and leveraged linear pushing actions along lines perpendicular to the gripper fingers to bring object centroids closer, thus enabled simultaneous grasping of objects initially outside the gripper’s opening width. Kishore et al. [20] proposed a method for efficiently transporting cups, bowls, and utensils on tables by combining pushing and stacking actions for simultaneous grasping. The method especially employed inward-pushing actions to reduce positional uncertainty.
The pushing-based methods have addressed the problem of scattered objects to some extent. However, they are predominantly rule-based, making it challenging to generate complex pushing trajectories required for handling intricate grouping and grasping tasks. In this paper, we propose employing imitation learning to generate complex pushing trajectories and solve the constraints on combinations of objects that can be grasped simultaneously. Previously, reinforcement learning methods have attracted much attention in robotic pushing and grasping as they can adapt to various object arrangements. For instance, Zeng et al. [21] achieved efficient grasping by enabling the complementary relationship between pushing and grasping actions to be learned through self-supervised deep reinforcement learning. Chen et al. [22] proposed a grasping strategy that combines rule-based methods with reinforcement learning to integrate pushing and grasping actions. Wang et al. [23] attained high success rates and robustness by leveraging self-supervised deep reinforcement learning to learn an integrated model of pushing and grasping actions. These studies demonstrate that reinforcement learning has the advantage of adapting to unknown environments.
@@
A key challenge in reinforcement learning lies in the design of rewards. This issue becomes particularly pronounced in the context of this study as we use pushing to achieve grouping, followed by grasping. Designing rewards that facilitate rapid convergence for the pushing-grouping-grasping routine is notably difficult. A learning method conceptually similar to reinforcement learning but not dependent on reward design is imitation learning. Recent studies have showcased the potential of imitation learning in robotic manipulation applications. For example, ACT [24] has achieved high success rates in delicate tasks such as opening and closing Ziploc bags and inserting batteries into controllers. Diffusion policy [3] has enabled high success rates in tasks requiring diverse actions and significant flexibility, such as spreading sauce on pizza dough or relocating T-shape blocks and mugs placed at random positions on a plane to target positions and orientations. Given its advantages, this research adopts imitation learning as the primary methodology. By learning flexible pushing actions and grasping decisions from human demonstrations, our approach enables dynamic selection of object combinations, rearrangement through complex pushing trajectories, and adaptive grasping sequences that were previously difficult to attain with rule-based or reinforcement learning methods.
@@ III Data Collection Using Visual Teleoperation
Similar to AnyTeleop [26], we develop a teleoperation system for a robot by detecting the hand skeleton pose using an RGB-D camera. Specifically, we employ the WiLoR [27] model for detecting the hand skeleton pose. This model estimates the 3D positions of keypoints on the hand based on a single RGB image. The output is described in the $\Sigma_{\textnormal{wilor}}$ coordinate system shown in Fig. 2(a). The origin of the coordinate system is defined at the wrist position, with the z-axis aligned with the camera’s line of sight, the x-axis corresponding to the width direction of the captured image, and the y-axis representing the height direction. The keypoints, including the positions on the wrist, index finger, and thumb used in this method, are illustrated in Fig. 2(b). After extracting the 3D hand skeleton data within the above coordinate system, we further integrate the data with the wrist depth value obtained from the depth frame of the same RGB-D camera and derive the 3D hand skeleton data in the coordinate system of the depth camera, $\Sigma_{\textnormal{cam}}$. After that, we transform the 3D data into commands for the robot, which includes the robot gripper’s position, orientation, and open/close states, thereby enabling the teleoperation of the robot. The specific mapping between the coordinate system and the robot’s control is defined as follows.
@@
Figure 2: Mapping relationship between hand poses detected by WiLoR and the end-effector of a robot. Teleoperation is achieved through the mapping to control the robot and collect training data.
@@
Jaw bottom of the robot gripper: The jaw bottom center of the robot’s end-effector is defined as a position corresponding to the MCP joint of the index finger and is transformed to the world coordinate system using the following equation:
[$\prescript{\Sigma_{\text{world}}}{}{\begin{bmatrix}\boldsymbol{p}_{\text{eef}}\\ 1\end{bmatrix}}={}^{\Sigma_{\text{world}}}\textbf{T}_{\Sigma_{\text{cam}}}{}^{\Sigma_{\text{cam}}}\textbf{T}_{\Sigma_{\text{wilor}}}\prescript{\Sigma_{\text{wilor}}}{}{\begin{bmatrix}\boldsymbol{p}_{5}\\ 1\end{bmatrix}}$ (1)
@@
Rotation of the robot gripper: The rotation of the end-effector, ${}^{\Sigma_{\text{world}}}\boldsymbol{R}_{\textnormal{eef}}$, is constrained such that its approaching direction is always vertically downward (the last columne or $\boldsymbol{z}$ axis of ${}^{\Sigma_{\text{world}}}\boldsymbol{R}_{\textnormal{eef}}$ is parallel with and inversely aligns with the $\boldsymbol{z}$ axis of $\Sigma_{\textnormal{world}}$ in Fig. 2(a)). The rotation around the approaching direction is defined using the position of the thumb tip (${}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{4}$) and the position of the index finger tip (${}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{8}$) following equation (2).
${}^{\Sigma_{\text{world}}}\boldsymbol{R}_{\text{eef}}=\begin{bmatrix}\hat{\boldsymbol{d}}_{x}&\hat{\textbf{d}}_{y}&0\\ \hat{\boldsymbol{d}}_{y}&-\hat{\boldsymbol{d}}_{x}&0\\ 0&0&1\end{bmatrix},~{}\hat{\boldsymbol{d}}=\frac{{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{4}}{\|{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{4}\|}$ (2)
@@
Gripping state: The robot gripper’s open/close state, $g$, is defined based on the distance between the thumb and the index finger following.
$g=\begin{cases}0&\text{if }d\geq\mu_{\text{up}}\\ 1&\text{if }d<\mu_{\text{bot}}\end{cases},~{}d=\frac{\|{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{4}\|}{\max\|{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wilor}}}\boldsymbol{p}_{4}\|}$ (3)
@@
Based on these mapping relationships, we define the observation $\boldsymbol{o}_{t}$ at a timestamp $t$ as
$\boldsymbol{o}_{t}=\left\{\boldsymbol{i}^{\text{gcam}}_{t},\boldsymbol{i}^{\text{lcam}}_{t},\boldsymbol{s}_{t}\right\},$ (4)
where $\boldsymbol{i}^{\text{gcam}}_{t},\boldsymbol{i}^{\text{lcam}}_{t}$ are video frames (images) captured by the global and local cameras. $\boldsymbol{s}_{t}$ is the state of the robot and is defined as
$\boldsymbol{s}_{t}=\left\{{p}_{x},{p}_{y},{p}_{z},{q}_{w},{q}_{x},{q}_{y},{q}_{z},{g}\right\},$ (5)
where $({p}_{x},{p}_{y},{p}_{z})$ are the ${}^{\Sigma_{\text{world}}}\boldsymbol{p}_{\text{eef}}$ position of the robot at time $t$. $({q}_{w},{q}_{x},{q}_{y},{q}_{z})$ are the ${}^{\Sigma_{\text{world}}}\textbf{R}_{\text{eef}}$ rotation of the robot at time $t$. $g$ is the gripper open/close state at time $t$.
@@
It has the same mathematical representation as $\boldsymbol{s}_{t}$, except that the constituent values denotes the position, rotation, and gripper state that the robot is going to act to at the next times stamp.

### v3

# Learning to Push, Group, and Grasp: A Diffusion Policy Approach for Multi-Object Delivery

@@ Abstract
Simultaneously grasping and delivering multiple objects can significantly enhance robotic work efficiency and has been a key research focus for decades. The primary challenge lies in determining how to push objects, group them, and execute simultaneous grasping for respective groups while considering object distribution and the hardware constraints of the robot. Traditional rule-based methods struggle to flexibly adapt to diverse scenarios. To address this challenge, this paper proposes an imitation learning-based approach. We collect a series of expert demonstrations through teleoperation and train a diffusion policy network, enabling the robot to dynamically generate action sequences for pushing, grouping, and grasping, thereby facilitating efficient multi-object grasping and delivery. We conducted experiments to evaluate the method under different training dataset sizes, varying object quantities, and real-world object scenarios. The results demonstrate that the proposed approach can effectively and adaptively generate multi-object grouping and grasping strategies. With the support of more training data, imitation learning is expected to be an effective approach for solving the multi-object grasping problem.
@@
Multi-Object Grasping, Imitation Learning
@@ I Introduction
This study focuses on enabling robots to perform simultaneous multi-object grasping while considering pushing-based grouping. Previous studies on simultaneous multi-object grasping have identified several challenges. For example, Sakamoto et al. [1] proposed a rule-based method to push objects closer together before grasping them simultaneously. However, the method cannot handle complex pushing actions, and the combinations of objects that can be grasped together are limited. Agboh et al. [2] studied identifying near objects arranged on a plane and grasping multiple objects by pushing them together as the gripper closes. While effective under specific conditions, the method assumes that objects are initially spaced within the gripper’s maximum opening width. Or else, it degenerates into single-object grasping. These limitations highlight the need for more advanced strategies that can group objects through flexible pushing actions and select suitable object combinations for simultaneous grasping. Designing such strategies through hand-crafted rules, however, remains a challenging task.
@@
Figure 1: (a) Collecting human expert data using teleoperation. (b) The collected dataset is used to train a diffusion policy network for inferring actions. (c) The policy network uses past observations to generate pushing or grasping actions.
@@
In this context, we propose addressing these issues by using imitation learning and enabling robots to learn how humans decide which objects to group and grasp. By acquiring grasping strategies through imitation learning, the proposed method enables robots to automatically choose suitable object combinations and perform the necessary pushing actions to group them. The method can overcome the limitations of existing methods and improve the efficiency and versatility of robotic picking tasks. Fig. 1 illustrates the components and workflow of the proposed framework. In the data collection phase shown in Fig. 1(a), a teleoperation system is developed for human operators to control the robot using natural hand movements. The setup enables collecting data like robot joint angles, gripper opening and closing states, and images from the global and local cameras mounted in the environment and on the robot hand. In the learning phase shown in Fig. 1(b), the framework employs Diffusion Policy [3] to learn appropriate robot actions several steps ahead, based on the robot’s state from previous steps. Finally, in the inference phase shown in Fig. 1(c), the framework uses the robot’s gripper poses, opening states, and camera images observed from a few steps before to predict new gripper poses and states for subsequent steps, thereby generating pushing and simultaneous grasping motions for multiple objects.
@@
Four groups of experiments were conducted to validate the proposed framework. The first group focused on comparing and analyzing the influence of different training data volumes and object quantities on learning and inference. The second group conducted a comparative study against a rule-based analytical method. The third group conducted fine-tuning of models pre-trained on a dataset of simple objects to more practical scenarios and evaluated the generality of the proposed method to real-world objects. The fourth group evaluates the generalization ability of the learned policy across different object geometries. The experimental results validated the effectiveness of the proposed method and the adaptivity of fine-tuned models. Additionally, the results indicated that training a well-performing policy network requires a large amount of high-quality data, with a positive correlation between data volume and learning performance. Developing efficient methods for collecting such data could potentially improve both learning and inference performance.
To the best of our knowledge, this is the first work to apply Diffusion Policy to choosing among pushing, grouping, and multi-object grasping. The proposed method is simple to implement and data-driven. We provide extensive experiments and in-depth analysis, and discuss both the strengths and limitations of the method. With larger-scale training data, imitation learning is expected to become an even more effective solution for multi-object grasping. We hope our findings and insights can inspire further research in this direction.
@@ II Related Work
Multi-object simultaneous grasping has been attracting significant attention for decades. As seminal work, Aiyama et al. [4] studied using two manipulators to cooperatively grasp and lift boxes. Harada and Kaneko [5][6] studied the neighboring equilibrium of multiple cylindrical objects and their enveloping grasps. Yoshikawa et al. [7] proposed the optimal power grasp condition for multiple objects from the viewpoint of reducing finger joint torques. Yamada et al. [8][9][10] analyzed the grasp stability of multiple objects by considering object surface curvatures and contact spring models.
@@
More recent studies by Chen et al. [11] inserted the Barrett hand into a pile of objects and estimated the number of grasped items using a deep learning model. Shenoy et al. [12] focused on efficiently transferring the grasped objects to a separate container. Takahashi et al. [13] studied grasping a given amount of granular foods. Post-grasping methods were explored for fine adjustments [14]. Li et al. [15] developed a framework for planning and learning multi-object grasping (lifting). Jiang et al. [16] developed a vacuum gripper equipped with multiple suction cups and the related grasp planning methods to simultaneously pick multiple objects from a tray. Nguyen et al. [17] proposed the Wiring-Claw Gripper, which took advantage of the soft interaction between wires in a claw and objects for simultaneous multi-object grasping. Yao et al. [18] used a dexterous hand with progressive planning to exploit redundancy for multi-object grasping.
@@
The above approaches could achieve stable grasping of multiple or unknown objects. However, they often assumed that the objects were initially positioned in proximity and were less applicable to scattered scenes. To tackle such limitations, several studies have explored using pushing actions to rearrange objects before simultaneous grasping. For example, Sakamoto et al. [1] proposed a method that can determine whether to grasp a single object, grasp two objects simultaneously, or push one object closer for simultaneous grasping based on the distance and friction between objects. Agboh et al. [2][19] proposed the $\mu$-MOG framework to simultaneously grasp multiple rigid convex polygonal objects scattered on a plane. The method took advantage of a gripper’s opening jaw to align objects before grasping. Shrey et al. [20] leveraged linear pushing actions along lines perpendicular to the gripper fingers to bring object centroids closer, allowing a robot to grasp multiple objects that initially lay beyond the gripper’s opening width. Ye et al. [21] used weighted graphs to model object distributions and assess collision-free clusters, thereby enabling efficient and reliable multi-object picking. Kishore et al. [22] proposed a method for efficiently delivering cups, bowls, and utensils on tables by combining pushing and stacking actions. The method especially employed inward-pushing to reduce positional uncertainty.
@@
The pushing-based methods have addressed the problem of scattered objects to some extent. However, they are predominantly rule-based, making it challenging to generate complex pushing trajectories for intricate grouping and grasping. Previously, reinforcement learning methods have attracted much attention in robotic pushing and grasping as they can adapt to various object arrangements. For instance, Zeng et al. [23] achieved efficient grasping by enabling the complementary relationship between pushing and grasping actions to be learned through self-supervised deep reinforcement learning. Chen et al. [24] proposed a grasping strategy that combines rule-based methods with reinforcement learning to integrate pushing and grasping. Wang et al. [25] attained high success rates and robustness by leveraging self-supervised deep reinforcement learning to learn an integrated model of pushing and grasping. These studies demonstrate that reinforcement learning has the advantage of adapting to unknown environments. However, a key challenge in reinforcement learning lies in the design of rewards. This issue becomes particularly pronounced in the context of this study as we use pushing to achieve grouping, followed by grasping. Designing rewards that facilitate rapid convergence for the pushing-grouping-grasping routine is notably difficult. A learning method conceptually similar to reinforcement learning but not dependent on reward design is imitation learning. Recent studies have showcased the potential of imitation learning in robotic manipulation applications. For example, ACT [26] has achieved high success rates in delicate tasks such as opening and closing Ziploc bags and inserting batteries into controllers. Diffusion policy [3] has enabled high success rates in tasks requiring diverse actions and significant flexibility, such as spreading sauce on pizza dough or relocating T-shape blocks and mugs placed at random positions on a plane to target positions and orientations. Given its advantages, this research adopts imitation learning as the primary methodology. By learning flexible pushing actions and grasping decisions from human demonstrations, our approach enables dynamic selection of object combinations, rearrangement through complex pushing trajectories, and adaptive grasping sequences that were previously difficult to attain with rule-based or reinforcement learning methods.
@@ III Data Collection Using Visual Teleoperation
Similar to AnyTeleop [27], we develop a teleoperation system for a robot by detecting the hand skeleton pose using an RGB-D camera. We use WiLoR [28] for detecting the hand skeleton pose. The output of WiLoR is a 3D skeleton described in the $\Sigma_{\text{wl}}$ coordinate system shown in Fig. 2. It is then integrated with depth values of the same RGB-D camera to derive the 3D skeleton in the coordinate system of the depth camera, $\Sigma_{\textnormal{cam}}$. After that, we transform the 3D data into actions for the robot, thereby enabling teleoperation. The coordinates and transformations used in the process are as follows.
@@
Figure 2: Mapping relationship between hand poses detected by WiLoR and the end-effector of the target robot.
@@
Jaw bottom of the robot gripper: The jaw bottom center of the robot’s end-effector is defined as a position corresponding to the MCP joint of the index finger ($\boldsymbol{p}_{5}$ in Fig. 2) and is transformed to the world coordinate system using
${}^{\Sigma_{\text{wd}}}\boldsymbol{p}_{\text{eef}}={}^{\Sigma_{\text{wd}}}\mathbf{R}_{\Sigma_{\text{cam}}}\cdot\left({}^{\Sigma_{\text{cam}}}\mathbf{R}_{\Sigma_{\text{wl}}}\cdot{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{5}+{}^{\Sigma_{\text{cam}}}\mathbf{p}_{\text{wl}}\right)+{}^{\Sigma_{\text{wd}}}\mathbf{p}_{\text{cam}}$
@@
Rotation of the robot gripper: The rotation of the end-effector, ${}^{\Sigma_{\text{wd}}}\boldsymbol{R}_{\textnormal{eef}}$, is constrained such that its approaching direction is always vertically downward (the last column or $\boldsymbol{z}$ axis of ${}^{\Sigma_{\text{wd}}}\boldsymbol{R}_{\textnormal{eef}}$ is parallel with and inversely aligns with the $\boldsymbol{z}$ axis of $\Sigma_{\textnormal{world}}$ in Fig. 2(a)). The rotation around the approaching direction is defined using the position of the thumb tip (${}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{4}$) and the position of the index finger tip (${}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{8}$) following equation (III).
${}^{\Sigma_{\text{wd}}}\boldsymbol{R}_{\text{eef}}=\begin{bmatrix}\hat{\boldsymbol{d}}_{x}&\hat{\boldsymbol{d}}_{y}&0\\ \hat{\boldsymbol{d}}_{y}&-\hat{\boldsymbol{d}}_{x}&0\\ 0&0&1\end{bmatrix},~\hat{\boldsymbol{d}}=\frac{{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{4}}{\|{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{4}\|}$
@@
Gripping state: The robot gripper’s open/close state, $g$, is defined based on the distance between the thumb and the index finger following
$g=\begin{cases}0&\text{if }d\geq\mu_{\text{up}}\\ 1&\text{if }d<\mu_{\text{bot}}\end{cases},~d=\frac{\|{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{4}\|}{\max\|{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{8}-{}^{\Sigma_{\text{wl}}}\boldsymbol{p}_{4}\|}$
@@
Based on these mapping relationships, we define the observation $\boldsymbol{o}_{t}$ at a timestamp $t$ as $\boldsymbol{o}_{t}=\{\boldsymbol{i}^{\text{gcam}}_{t},\boldsymbol{i}^{\text{lcam}}_{t},\boldsymbol{s}_{t}\}$, where $\boldsymbol{i}^{\text{gcam}}_{t},\boldsymbol{i}^{\text{lcam}}_{t}$ are video frames captured by the global and local cameras. $\boldsymbol{s}_{t}$ is the state of the robot and is defined as
$\boldsymbol{s}_{t}=\left\{{p}_{x},{p}_{y},{p}_{z},{q}_{w},{q}_{x},{q}_{y},{q}_{z},{g}\right\}.$
Here, $({p}_{x},{p}_{y},{p}_{z})$ are the ${}^{\Sigma_{\text{wd}}}\boldsymbol{p}_{\text{eef}}$ position of the robot at time $t$. $({q}_{w},{q}_{x},{q}_{y},{q}_{z})$ are the ${}^{\Sigma_{\text{wd}}}\textbf{R}_{\text{eef}}$ rotation of the robot at time $t$. $g$ is the gripper open/close state at time $t$.
@@
It has the same mathematical representation as $\boldsymbol{s}_{t}$, except that the constituent values denote the position, rotation, and gripper state that the robot is going to act on at the next time step.We clarified the above discussion at the beginning of the experimental section.
During data collection, a human operator controls the robot via the teleoperation system to demonstrate how to group and grasp multiple objects. The collected demonstrations were used to train a Diffusion Policy to learn human strategies. Further details are provided in the following section Note 1 Note 1 Note 1A follow-up interview with the human operator revealed the overall heuristics used during teleoperated data collection. To avoid overloading the gripper, the number of grouped objects was limited to three, based on object size. Objects were grouped if they were spatially close and roughly aligned to avoid significant rotational adjustments. When choosing the first object, the operator tended to select one surrounded by nearby candidates to facilitate multi-object grouping. These human strategies may offer useful inspiration for readers interested in understanding the underlying rules..